# Generalized open-loop Nash equilibria in linear-quadratic difference games with coupled-affine inequality constraints

### v5

@@ 1 Introduction
We denote the sets of natural numbers, real numbers, $n$-dimensional Euclidean space, $n$-dimensional non-negative orthant, and $n\times m$ real matrices by $\mathbb{N}$, $\mathbb{R}$, $\mathbb{R}^{n}$, +$\mathbb{R}^{n}_{+}$, and $\mathbb{R}^{n\times m}$, respectively. The transposes of a vector $a$ and a matrix $A$ are denoted by $a^{\prime}$ and $A^{\prime}$. The quadratic form $a^{\prime}Aa$ is denoted by $\|a\|_{A}$. For $A\in\mathbb{R}^{n\times n}$ and $a\in\mathbb{R}^{n}$, where $n=n_{1}+\cdots+n_{K}$, $[A]_{ij}$ denotes the $n_{i}\times n_{j}$ submatrix, and $[a]_{i}$ denotes the $n_{i}\times 1$ subvector. Column vectors $[v_{1}^{\prime},\cdots,v_{n}^{\prime}]^{\prime}$ are written as $\mathrm{col}\{v_{1},\cdots,v_{n}\}$ or $\mathrm{col}\{v_{k}\}_{k=1}^{n}$. The identity and zero matrices of appropriate dimensions are denoted by $\mathbf{I}$ and $\mathbf{0}$, respectively. The block diagonal matrix with diagonal elements $M_{1},\cdots,M_{K}$ is denoted by $\oplus_{k=1}^{K}M_{k}$. The Kronecker product is $\otimes$. Vectors $x,y\in\mathbb{R}^{n}$ are complementary if $x\geq 0$, $y\geq 0$, and $x^{\prime}y=0$, denoted by $0\leq x\perp y\geq 0$.
@@ 2 Dynamic game with coupled constraints
where $M_{k}^{i}\in\mathbb{R}^{c_{i}\times n},N_{k}^{i}\in\mathbb{R}^{c_{i}\times m},r_{k}^{i}\in\mathbb{R}^{c_{i}}$. For player $i\in\mathsf{N}$ we denote $-i:=\mathsf{N}\setminus\{i\}$. At any instant $k\in\mathsf{K}_{l}$ the collection of actions of all players except player $i$ be denoted by $u^{-i}_{k}:=\mathrm{col}\{u_{k}^{1},\cdots,u_{k}^{i-1},u_{k}^{i+1},\cdots,u_{k}^{N}\}$. The profile of actions, also referred to as a strategy, of player $i\in\mathsf{N}$ be denoted by $\mathsf{u}^{i}:=\mathrm{col}\{u_{k}^{i}\}_{k=0}^{K-1}$, and the strategies of all players except player $i$ be denoted by $\mathsf{u}^{-i}:=\mathrm{col}\{u_{k}^{-i}\}_{k=0}^{K-1}$. Each player $i\in\mathsf{N}$, using his strategy $\mathsf{u}^{i}$, seeks to minimize the following interdependent stage-additive cost functional
@@ 3 Generalized Open-loop Nash Equilibrium
[…]
@@ 3.2.2 Main result
Let Assumptions 1 and 2 hold. Let the matrices $\{Y_{k}^{i},~{}k\in\mathsf{K}_{l},~{}i\in\mathsf{N}\}$, as defined in Theorem 1, be positive-definite. Then, SGCC is a convex game and a GOLNE exists for SGCC.
@@
DGCLCS1SGCCLCS2$\{\bm{x}^{\dagger},\bm{\lambda},\bm{\mu}^{\dagger}\}^{1}$$\{\bm{x}^{\dagger},\bm{e},\bm{\mu}^{\dagger};\bm{b}\}^{2}_{\bm{\mu}^{\dagger}}$$\{Y_{k}^{i}\succ 0,~{}k\in\mathsf{K}_{l},~{}i\in\mathsf{N}\}$Theorem 1N.C.N.C.Theorem 2Theorem 3Theorem 4S.C.
Figure 1: N.C. (S.C.) refer to necessary (sufficient) conditions. Replace $\dagger$ with $\star$ for Theorem 2 and $\dagger$ with -- for Theorem 3.
@@ 3.3 Reformulation of LCS1 as a linear complementarity problem
×$\quad\times\sum_{j\in\mathsf{N}}B^{j}_{k}(R^{jj}_{k})^{-1}({B^{j}_{k}}^{\prime}\zeta^{j}_{k+1}-[N^{j}_{k}]_{j}^{\prime}\mu_{k}^{j\star})+p_{k}^{i},~{}\zeta_{K}^{i}=p_{K}^{i}.$ (33b)
@@
$0\leq(\overline{\mathbf{M}}_{k}+\overline{\mathbf{N}}_{k}\mathbf{F}_{k})x_{k}^{\star}+\overline{\mathbf{N}}_{k}\overline{\mathbf{F}}_{k+1}\bm{\zeta}_{k+1}+\overline{\mathbf{N}}_{k}\tilde{\mathbf{F}}_{k}\bm{\upmu}_{k}^{\star}+\mathbf{r}_{k}\perp\bm{\upmu}_{k}^{\star}\geq 0$ (36d)
@@
$0\leq\mathsf{M}_{\mathsf{K}}x_{\mathsf{K}}^{\star}+\mathsf{N}_{\mathsf{K}}\bm{\zeta}_{\mathsf{K}}+\tilde{\mathsf{N}}_{\mathsf{K}}\bm{\upmu}_{\mathsf{K}}^{\star}+\mathbf{r}_{\mathsf{K}}\perp\bm{\upmu}_{\mathsf{K}}^{\star}\geq 0$ (39d)
@@ 4 Numerical Illustration
[…]
@@ 5 Conclusions
We define $\mathbf{R}_{k}=\oplus_{i=1}^{N}R_{k}^{ii}$, $\mathbf{B}_{k}=\oplus_{i=1}^{N}B_{k}^{i}$, $\mathbf{N}_{k}=\oplus_{i=1}^{N}[N^{i}_{k}]_{i}$, $\mathbf{M}_{k}=\oplus_{i=1}^{N}M^{i}_{k}$, $\mathbf{r}_{k}=\mathrm{col}\{r_{k}^{i}\}_{i=1}^{N}$, $\mathsf{u}_{k}^{\star}=\mathrm{col}\{u_{k}^{i\star}\}_{i=1}^{N}$, $\bm{\upmu}_{k}^{\star}=\mathrm{col}\{\mu_{k}^{i\star}\}_{i=1}^{N}$,  $k\in\mathsf{K}_{l}$, $\mathbf{p}_{k}=\mathrm{col}\{p_{k}^{i}\}_{i=1}^{N}$, $P_{k}=\mathrm{col}\{P_{k}^{i}\}_{i=1}^{N}$, $\bm{\zeta}_{k}=\mathrm{col}\{\zeta_{k}^{i}\}_{i=1}^{N}$, $k\in\mathsf{K}$.   In (36): $\mathbf{G}_{k}=(\Lambda_{k})^{-1}A_{k}$, $[\overline{\mathbf{G}}_{k+1}]_{i}=-(\Lambda_{k})^{-1}B^{i}_{k}(R^{ii}_{k})^{-1}{B^{i}_{k}}^{\prime}$, $[\tilde{\mathbf{G}}_{k}]_{i}=(\Lambda_{k})^{-1}B^{i}_{k}(R^{ii}_{k})^{-1}[N^{i}_{k}]_{i}^{\prime}$, $\mathbf{F}_{k}=-(\mathbf{R}_{k})^{-1}\mathbf{B}_{k}^{\prime}P_{k+1}\mathbf{G}_{k}$, $\overline{\mathbf{F}}_{k+1}=-(\mathbf{R}_{k})^{-1}\mathbf{B}_{k}^{\prime}(\mathbf{I}+P_{k+1}\overline{\mathbf{G}}_{k+1})$, $\tilde{\mathbf{F}}_{k}=(\mathbf{R}_{k})^{-1}(\mathbf{N}_{k}^{\prime}-\mathbf{B}_{k}^{\prime}P_{k+1}\tilde{\mathbf{G}}_{k})$, $\mathbf{H}_{k+1}=(\mathbf{I}_{N}\otimes A_{k}^{\prime})(\mathbf{I}+P_{k+1}\overline{\mathbf{G}}_{k+1})$, $\overline{\mathbf{H}}_{k}=(\mathbf{I}_{N}\otimes A_{k}^{\prime})P_{k+1}\tilde{\mathbf{G}}_{k}-\mathbf{M}_{k}^{\prime}$ for $k\in\mathsf{K}_{l}$.  For the state transition matrices in (36b)-(36c): $\phi(k,\tau)=\mathbf{G}_{k-1}\mathbf{G}_{k-2}\cdots\mathbf{G}_{\tau}$ for $k>\tau$ and $\phi(k,\tau)=\mathbf{I}$ for $k=\tau$, $k,\tau\in\mathsf{K}$. $\varphi(k,\tau)=\mathbf{H}_{k+1}\mathbf{H}_{k+2}\cdots\mathbf{H}_{\tau}$ for $k<\tau$ and $\varphi(k,\tau)=\mathbf{I}$ for $k=\tau$, $k,\tau\in\mathsf{K}_{l}$.  In (39): $x_{\mathsf{K}}^{\star}=\mathrm{col}\{x_{k}^{\star}\}_{k=0}^{K-1}$, $\mathsf{u}^{\star}_{\mathsf{K}}=\mathrm{col}\{\mathsf{u}^{\star}_{k}\}_{k=0}^{K-1}$, $\bm{\upmu}^{\star}_{\mathsf{K}}=\mathrm{col}\{\bm{\upmu}^{\star}_{k}\}_{k=0}^{K-1}$, $\mathbf{p}_{\mathsf{K}}=\mathrm{col}\{\mathbf{p}_{k+1}\}_{k=0}^{K-1}$, $\bm{\zeta}_{\mathsf{K}}=\mathrm{col}\{\bm{\zeta}_{k+1}\}_{k=0}^{K-1}$, $\textbf{r}_{\mathsf{K}}=\mathrm{col}\{\textbf{r}_{k}\}_{k=0}^{K-1}$, $\mathbf{F}_{\mathsf{K}}=\oplus_{k=0}^{K-1}\mathbf{F}_{k}$, $\overline{\mathbf{F}}_{\mathsf{K}}=\oplus_{k=0}^{K-1}\overline{\mathbf{F}}_{k+1}$, $\tilde{\mathbf{F}}_{\mathsf{K}}=\oplus_{k=0}^{K-1}\tilde{\mathbf{F}}_{k}$, $[\bm{\Phi}_{0}]_{k}=\phi(k-1,0)$, $[\bm{\Phi}_{1}]_{1\tau}=\mathbf{0}$, $[\bm{\Phi}_{2}]_{1\tau}=\mathbf{0}$, $[\bm{\Phi}_{1}]_{k\tau}=\sum_{\rho=1}^{\mathrm{min}\,\,(k-1,\tau-1)}\phi(k-1,\rho)\overline{\mathbf{G}}_{\rho}\varphi(\rho,\tau-1)$ for $k>1$, $[\bm{\Phi}_{2}]_{k\tau}=\phi(k-1,1)\tilde{\mathbf{G}}_{0}$ for $k>\tau=1$, $[\bm{\Phi}_{2}]_{k\tau}=\big{(}\sum_{\rho=1}^{\tau-1}\phi(k-1,\rho)\overline{\mathbf{G}}_{\rho}\varphi(\rho,\tau-1)\big{)}\overline{\mathbf{H}}_{\tau-1}+\phi(k-1,\tau)\tilde{\mathbf{G}}_{\tau-1}$ for $1<\tau<k$, $[\bm{\Phi}_{2}]_{k\tau}=\big{(}\sum_{\rho=1}^{k-1}\phi(k-1,\rho)\overline{\mathbf{G}}_{\rho}\varphi(\rho,\tau-1)\big{)}\overline{\mathbf{H}}_{\tau-1}$ for $\tau\geq k$ $[\bm{\Psi}_{1}]_{k\tau}=\varphi(k,\tau)$, $[\bm{\Psi}_{2}]_{k\tau}=\varphi(k,\tau-1)\overline{\mathbf{H}}_{\tau-1}$ for $\tau\geq k$, $[\bm{\Psi}_{1}]_{k\tau}=\mathbf{0}$, $[\bm{\Psi}_{2}]_{k\tau}=\mathbf{0}$ for $\tau<k$ with $k,\tau\in\mathsf{K}_{r}$, $\mathsf{M}_{\mathsf{K}}=\oplus_{k=0}^{K-1}(\overline{\mathbf{M}}_{k}+\overline{\mathbf{N}}_{k}\mathbf{F}_{k})$, $\mathsf{N}_{\mathsf{K}}=\oplus_{k=0}^{K-1}(\overline{\mathbf{N}}_{k}\overline{\mathbf{F}}_{k+1})$ and $\tilde{\mathsf{N}}_{\mathsf{K}}=\oplus_{k=0}^{K-1}(\overline{\mathbf{N}}_{k}\tilde{\mathbf{F}}_{k})$.

### v6

@@ 1 Introduction
We denote the sets of natural numbers, real numbers, $n$-dimensional Euclidean space, $n$-dimensional non-negative orthant, and $n\times m$ real matrices by $\mathbb{N}$, $\mathbb{R}$, $\mathbb{R}^{n}$, $\mathbb{R}^{n}_{+}$, and $\mathbb{R}^{n\times m}$, respectively. The transposes of a vector $a$ and a matrix $A$ are denoted by $a^{\prime}$ and $A^{\prime}$. The quadratic form $a^{\prime}Aa$ is denoted by $\|a\|_{A}$. For $A\in\mathbb{R}^{n\times n}$ and $a\in\mathbb{R}^{n}$, where $n=n_{1}+\cdots+n_{K}$, $A]_{ij}$ denotes the $n_{i}\times n_{j}$ submatrix, and $a]_{i}$ denotes the $n_{i}\times 1$ subvector. Column vectors $v_{1}^{\prime},\cdots,v_{n}^{\prime}]^{\prime}$ are written as $\mathrm{col}\{v_{1},\cdots,v_{n}\}$ or $\mathrm{col}\{v_{k}\}_{k=1}^{n}$. The identity and zero matrices of appropriate dimensions are denoted by $\mathbf{I}$ and $\mathbf{0}$, respectively. The block diagonal matrix with diagonal elements $M_{1},\cdots,M_{K}$ is denoted by $\oplus_{k=1}^{K}M_{k}$. The Kronecker product is $\otimes$. Vectors $x,y\in\mathbb{R}^{n}$ are complementary if $x\geq 0$, $y\geq 0$, and $x^{\prime}y=0$, denoted by $0\leq x\perp y\geq 0$.
@@ 2 Dynamic game with coupled constraints
where $M_{k}^{i}\in\mathbb{R}^{c_{i}\times n},N_{k}^{i}\in\mathbb{R}^{c_{i}\times m},r_{k}^{i}\in\mathbb{R}^{c_{i}}$. For player $i\in\mathsf{N}$ we denote $i:=\mathsf{N}\setminus\{i\}$. At any instant $k\in\mathsf{K}_{l}$ the collection of actions of all players except player $i$ be denoted by $u^{-i}_{k}:=\mathrm{col}\{u_{k}^{1},\cdots,u_{k}^{i-1},u_{k}^{i+1},\cdots,u_{k}^{N}\}$. The profile of actions, also referred to as a strategy, of player $i\in\mathsf{N}$ be denoted by $\mathsf{u}^{i}:=\mathrm{col}\{u_{k}^{i}\}_{k=0}^{K-1}$, and the strategies of all players except player $i$ be denoted by $\mathsf{u}^{-i}:=\mathrm{col}\{u_{k}^{-i}\}_{k=0}^{K-1}$. Each player $i\in\mathsf{N}$, using his strategy $\mathsf{u}^{i}$, seeks to minimize the following interdependent stage-additive cost functional
@@ 3 Generalized Open-loop Nash Equilibrium
$R(x_{0}):=\{$ $(\mathsf{u}^{i},\mathsf{u}^{-i})\in\mathbb{R}^{Km}~|~(x_{0},(\mathsf{u}^{i},\mathsf{u}^{-i}))\in\Omega\}.$ (5)
@@ 3.2.2 Main result
Let Assumptions 1 and 2 hold. Let the matrices $\{Y_{k}^{i},~k\in\mathsf{K}_{l},~i\in\mathsf{N}\}$, as defined in Assumption 2, be positive-definite. Then, SGCC is a convex game and a GOLNE exists for SGCC.
@@
[…]
Figure 1: N.C. (S.C.) refer to necessary (sufficient) conditions. Replace $\dagger$ with $\star$ for Theorem 2 and $\dagger$ with - for Theorem 3.
@@ 3.3 Reformulation of LCS1 as a linear complementarity problem
$\quad\times\sum_{j\in\mathsf{N}}B^{j}_{k}(R^{jj}_{k})^{-1}({B^{j}_{k}}^{\prime}\zeta^{j}_{k+1}-[N^{j}_{k}]_{j}^{\prime}\mu_{k}^{j\star})+p_{k}^{i},~\zeta_{K}^{i}=p_{K}^{i}.$ (33b)
@@
$0\leq(\overline{\mathbf{M}}_{k}+\overline{\mathbf{N}}_{k}\mathbf{F}_{k})x_{k}^{\star}+\overline{\mathbf{N}}_{k}\overline{\mathbf{F}}_{k+1}\bm{\zeta}_{k+1}+\overline{\mathbf{N}}_{k}\tilde{\mathbf{F}}_{k}\bm{\upmu}_{k}^{\star}+\mathbf{r}_{k}\perp\bm{\upmu}_{k}^{\star}\geq 0.$ (36d)
@@
$0\leq\mathsf{M}_{\mathsf{K}}x_{\mathsf{K}}^{\star}+\mathsf{N}_{\mathsf{K}}\bm{\zeta}_{\mathsf{K}}+\tilde{\mathsf{N}}_{\mathsf{K}}\bm{\upmu}_{\mathsf{K}}^{\star}+\mathbf{r}_{\mathsf{K}}\perp\bm{\upmu}_{\mathsf{K}}^{\star}\geq 0.$ (39d)
@@ 4 Numerical Illustration
$\text{Relay: }L_{k}^{l}=v_{k}^{1l}+v_{k}^{2l}\leq c^{l}>0,~l=1,2,$ (44a)
$\text{Destination: }\tilde{L}_{k}^{i}=v_{k}^{i1}+v_{k}^{i2}\leq\tilde{c}^{i}>0,~i=1,2,$ (44b)
$\text{Battery: }x_{k+1}^{l}=x_{k}^{l}-\delta^{l}(v_{k}^{1l}+v_{k}^{2l})\geq{b^{l}_{min}},~l=1,2,$ (44c)
$\text{Transmission: }0\leq v_{k}^{il}\leq\bar{v}^{il}=\tfrac{w^{il}}{t^{il}},~i,l=1,2.$ (44d)
@@ 5 Conclusions
We define $\mathbf{R}_{k}=\oplus_{i=1}^{N}R_{k}^{ii}$, $\mathbf{B}_{k}=\oplus_{i=1}^{N}B_{k}^{i}$, $\mathbf{N}_{k}=\oplus_{i=1}^{N}[N^{i}_{k}]_{i}$, $\mathbf{M}_{k}=\oplus_{i=1}^{N}M^{i}_{k}$, $\mathbf{r}_{k}=\mathrm{col}\{r_{k}^{i}\}_{i=1}^{N}$, $\mathsf{u}_{k}^{\star}=\mathrm{col}\{u_{k}^{i\star}\}_{i=1}^{N}$, $\bm{\upmu}_{k}^{\star}=\mathrm{col}\{\mu_{k}^{i\star}\}_{i=1}^{N}$,  $k\in\mathsf{K}_{l}$, $\mathbf{p}_{k}=\mathrm{col}\{p_{k}^{i}\}_{i=1}^{N}$, $P_{k}=\mathrm{col}\{P_{k}^{i}\}_{i=1}^{N}$, $\bm{\zeta}_{k}=\mathrm{col}\{\zeta_{k}^{i}\}_{i=1}^{N}$, $k\in\mathsf{K}$.   In (36): $\mathbf{G}_{k}=(\Lambda_{k})^{-1}A_{k}$, $\overline{\mathbf{G}}_{k+1}]_{i}=-(\Lambda_{k})^{-1}B^{i}_{k}(R^{ii}_{k})^{-1}{B^{i}_{k}}^{\prime}$, $\tilde{\mathbf{G}}_{k}]_{i}=(\Lambda_{k})^{-1}B^{i}_{k}(R^{ii}_{k})^{-1}[N^{i}_{k}]_{i}^{\prime}$, $\mathbf{F}_{k}=-(\mathbf{R}_{k})^{-1}\mathbf{B}_{k}^{\prime}P_{k+1}\mathbf{G}_{k}$, $\overline{\mathbf{F}}_{k+1}=-(\mathbf{R}_{k})^{-1}\mathbf{B}_{k}^{\prime}(\mathbf{I}+P_{k+1}\overline{\mathbf{G}}_{k+1})$, $\tilde{\mathbf{F}}_{k}=(\mathbf{R}_{k})^{-1}(\mathbf{N}_{k}^{\prime}-\mathbf{B}_{k}^{\prime}P_{k+1}\tilde{\mathbf{G}}_{k})$, $\mathbf{H}_{k+1}=(\mathbf{I}_{N}\otimes A_{k}^{\prime})(\mathbf{I}+P_{k+1}\overline{\mathbf{G}}_{k+1})$, $\overline{\mathbf{H}}_{k}=(\mathbf{I}_{N}\otimes A_{k}^{\prime})P_{k+1}\tilde{\mathbf{G}}_{k}-\mathbf{M}_{k}^{\prime}$ for $k\in\mathsf{K}_{l}$.  For the state transition matrices in (36b)-(36c): $\phi(k,\tau)=\mathbf{G}_{k-1}\mathbf{G}_{k-2}\cdots\mathbf{G}_{\tau}$ for $k>\tau$ and $\phi(k,\tau)=\mathbf{I}$ for $k=\tau$, $k,\tau\in\mathsf{K}$. $\varphi(k,\tau)=\mathbf{H}_{k+1}\mathbf{H}_{k+2}\cdots\mathbf{H}_{\tau}$ for $k<\tau$ and $\varphi(k,\tau)=\mathbf{I}$ for $k=\tau$, $k,\tau\in\mathsf{K}_{l}$.  In (39): $x_{\mathsf{K}}^{\star}=\mathrm{col}\{x_{k}^{\star}\}_{k=0}^{K-1}$, $\mathsf{u}^{\star}_{\mathsf{K}}=\mathrm{col}\{\mathsf{u}^{\star}_{k}\}_{k=0}^{K-1}$, $\bm{\upmu}^{\star}_{\mathsf{K}}=\mathrm{col}\{\bm{\upmu}^{\star}_{k}\}_{k=0}^{K-1}$, $\mathbf{p}_{\mathsf{K}}=\mathrm{col}\{\mathbf{p}_{k+1}\}_{k=0}^{K-1}$, $\bm{\zeta}_{\mathsf{K}}=\mathrm{col}\{\bm{\zeta}_{k+1}\}_{k=0}^{K-1}$, $\textbf{r}_{\mathsf{K}}=\mathrm{col}\{\textbf{r}_{k}\}_{k=0}^{K-1}$, $\mathbf{F}_{\mathsf{K}}=\oplus_{k=0}^{K-1}\mathbf{F}_{k}$, $\overline{\mathbf{F}}_{\mathsf{K}}=\oplus_{k=0}^{K-1}\overline{\mathbf{F}}_{k+1}$, $\tilde{\mathbf{F}}_{\mathsf{K}}=\oplus_{k=0}^{K-1}\tilde{\mathbf{F}}_{k}$, $\bm{\Phi}_{0}]_{k}=\phi(k-1,0)$, $\bm{\Phi}_{1}]_{1\tau}=\mathbf{0}$, $\bm{\Phi}_{2}]_{1\tau}=\mathbf{0}$, $\bm{\Phi}_{1}]_{k\tau}=\sum_{\rho=1}^{\mathrm{min}\,\,(k-1,\tau-1)}\phi(k-1,\rho)\overline{\mathbf{G}}_{\rho}\varphi(\rho,\tau-1)$ for $k>1$, $\bm{\Phi}_{2}]_{k\tau}=\phi(k-1,1)\tilde{\mathbf{G}}_{0}$ for $k>\tau=1$, $\bm{\Phi}_{2}]_{k\tau}=\big(\sum_{\rho=1}^{\tau-1}\phi(k-1,\rho)\overline{\mathbf{G}}_{\rho}\varphi(\rho,\tau-1)\big)\overline{\mathbf{H}}_{\tau-1}+\phi(k-1,\tau)\tilde{\mathbf{G}}_{\tau-1}$ for $1<\tau<k$, $\bm{\Phi}_{2}]_{k\tau}=\big(\sum_{\rho=1}^{k-1}\phi(k-1,\rho)\overline{\mathbf{G}}_{\rho}\varphi(\rho,\tau-1)\big)\overline{\mathbf{H}}_{\tau-1}$ for $\tau\geq k$ $\bm{\Psi}_{1}]_{k\tau}=\varphi(k,\tau)$, $\bm{\Psi}_{2}]_{k\tau}=\varphi(k,\tau-1)\overline{\mathbf{H}}_{\tau-1}$ for $\tau\geq k$, $\bm{\Psi}_{1}]_{k\tau}=\mathbf{0}$, $\bm{\Psi}_{2}]_{k\tau}=\mathbf{0}$ for $\tau<k$ with $k,\tau\in\mathsf{K}_{r}$, $\mathsf{M}_{\mathsf{K}}=\oplus_{k=0}^{K-1}(\overline{\mathbf{M}}_{k}+\overline{\mathbf{N}}_{k}\mathbf{F}_{k})$, $\mathsf{N}_{\mathsf{K}}=\oplus_{k=0}^{K-1}(\overline{\mathbf{N}}_{k}\overline{\mathbf{F}}_{k+1})$ and $\tilde{\mathsf{N}}_{\mathsf{K}}=\oplus_{k=0}^{K-1}(\overline{\mathbf{N}}_{k}\tilde{\mathbf{F}}_{k})$.